Truck driver at the border tool station to be cleared at a rate of two per hour the distribution of arrivals approximates a poisson distribution

What is the probability that no truck arrives in a particular hour

What is the probability that at least one truck arrives during a particular hour


1
Expert's answer
2021-06-27T03:56:47-0400

Let "X=" the number of tracks arrived: "X\sim Po(\lambda)."

Given "\lambda=2"


"P(X=0)=\dfrac{e^{-2}\cdot2^0}{0!}=e^{-2}\approx0.135335"

"P(X\geq 1)=1-P(X=0)=1-\dfrac{e^{-2}\cdot2^0}{0!}"

"=1-e^{-2}\approx0.864665"